OpenStudy (anonymous):

Which point lies on the line described by the equation below? y + 3 = 2(x - 1)

OpenStudy (pinksapphire):

You have to solve for y first.

OpenStudy (anonymous):

how?

OpenStudy (mrnood):

fist expand the bracket on the right of the equation

OpenStudy (pinksapphire):

y+3=2(x-1) -3 First you have to move the 3 to get y by itself. y=2(x-1)-3 This is your equation.

OpenStudy (anonymous):

oh

OpenStudy (pinksapphire):

Do you have a scientific calculator?

OpenStudy (anonymous):

yes

OpenStudy (pinksapphire):

Okay, go to y=

OpenStudy (anonymous):

ok

OpenStudy (pinksapphire):

Then, type the equation I just showed you.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank you

OpenStudy (pinksapphire):

Maybe I did something wrong in the equation...

OpenStudy (pinksapphire):

But if you go to 2nd graph, then you will see a table with coordinate. Pick a coordinate and substitute the coordinate in the equation and see if it works...

OpenStudy (pinksapphire):

What are you learning right now?

OpenStudy (pinksapphire):

Something is wrong in this equation...